OpenStudy (anonymous):

If I have 590 mL of a gas at -55 degrees Celsius, what would the volume be a 30 degrees Celsius? PLEASE show all steps taken so I know how to do this later...

OpenStudy (aaronq):

use this: \[\frac{ P _{1}V _{1} }{ T _{1} }=\frac{ P _{2}V _{2} }{ T _{2} }\]

OpenStudy (aaronq):

well i guess you can disregard the pressure, since theres no information about that, so I'm a asumming it's constant

OpenStudy (anonymous):

So it would be Volume/temprature=volume 2/temprature 2? As in 590/-55 equals V2/30? Then I would find V2? Then divide V2 by 30 to get the answer, right?

OpenStudy (aaronq):

nope, treat it like algebra, 590/-55=x/30 solve for x

OpenStudy (anonymous):

So the new volume would be -321.9 mL of gas?

OpenStudy (aaronq):

well i would convert the T to kelvin so you don't have a -ve temp

OpenStudy (anonymous):

I will. Thank you very much.